OpenStudy (anonymous):

Two equations are given below: a − 3b = 16 a = b − 2 What is the solution to the set of equations in the form (a, b)?

OpenStudy (kidrah69):

well substitute \[(b-2) − 3b = 16 \] it already tells u what a equals

OpenStudy (kidrah69):

solve that.....

OpenStudy (anonymous):

(−2, −6) (−7, −9) (−11, −9) (−12, −10)

OpenStudy (anonymous):

what is the ???

OpenStudy (kidrah69):

refresh

OpenStudy (anonymous):

um is it 1b=16?

OpenStudy (anonymous):

wait so then b=16? what?

OpenStudy (kidrah69):

incorrect \[b-2-3b=16\] whats b-3b?

OpenStudy (anonymous):

3b?

OpenStudy (anonymous):

i dont know this

OpenStudy (kidrah69):

1-3?

OpenStudy (anonymous):

2

OpenStudy (kidrah69):

no 1 MINUS 3

OpenStudy (kidrah69):

not 3-1

OpenStudy (anonymous):

-2* sorry

OpenStudy (kidrah69):

its oky :) since we have what is 16+2?

OpenStudy (anonymous):

18

OpenStudy (kidrah69):

whats -2/18?

OpenStudy (anonymous):

-0.11

OpenStudy (kidrah69):

18/-2

OpenStudy (anonymous):

-9

OpenStudy (kidrah69):

\[(b-2)-3b=16 \] \[b-2-3b=16\] subtract b \[b-3b=-2b\] you have -2b \[-2b-2=16\] add 2 to both sides to isolate b \[-2b=18\] divide to get b value b=-9 plug it in the other equation \[a=b-2 \] plug -9 \[a=-9-2\] whats -9-2? a=?

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

am i correct?

OpenStudy (kidrah69):

yep :)

OpenStudy (anonymous):

so then whats the answer? ? (−2, −6) (−7, −9) (−11, −9) (−12, −10)

OpenStudy (kidrah69):

u tell me if b=-9 a=-11 what is the answer?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

the 3rd one

OpenStudy (kidrah69):

good job :)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (kidrah69):

you're welcome :)